Packaging Solutions Corporation manufactures and sells a wide variety of packaging products. Performance reports are prepared monthly for each department. The planning budget and flexible budget for the Production Department are based on the following formulas, where q is the number of labor-hours worked in a month:

Cost Formulas
Direct labor $16.30q
Indirect labor $4,600 + $1.50q
Utilities $5,500 + $0.80q
Supplies $1,300 + $0.10q
Equipment depreciation $18,300 + $2.80q
Factory rent $8,200
Property taxes $2,700
Factory administration $13,300 + $0.90q

The Production Department planned to work 4,100 labor-hours in March; however, it actually worked 3,900 labor-hours during the month. Its actual costs incurred in March are listed below:

Actual Cost Incurred in March
Direct labor $ 65,150
Indirect labor $ 9,930
Utilities $ 9,170
Supplies $ 1,920
Equipment depreciation $ 29,220
Factory rent $ 8,600
Property taxes $ 2,700
Factory administration $ 16,220

Required:

1. Prepare the Production Department’s planning budget for the month.

2. Prepare the Production Department’s flexible budget for the month.

3. Prepare the Production Department’s flexible budget performance report for March, including both the spending and activity variances.
Answer #1
Cost Formulas Cost For q=4100 Cost for q=3900
Direct labor $16.30q $66,830 $63,570
Indirect labor $4,600 + $1.50q $10,750 $10,450
Utilities $5,500 + $0.80q $8,780 $8,620
Supplies $1,300 + $0.10q $1,710 $1,690
Equipment depreciation $18,300 + $2.80q $29,780 $29,220
Factory rent $8,200 $8,200 $8,200
Property taxes $2,700 $2,700 $2,700
Factory administration $13,300 + $0.90q $16,990 $16,810
1 PLANNING BUDGET
Budgeted Cost
Direct labor $66,830
Indirect labor $10,750
Utilities $8,780
Supplies $1,710
Equipment depreciation $29,780
Factory rent $8,200
Property taxes $2,700
Factory administration $16,990
Total $145,740
2 FLEXIBLE BUDGET
Budgeted Cost
Direct labor $63,570
Indirect labor $10,450
Utilities $8,620
Supplies $1,690
Equipment depreciation $29,220
Factory rent $8,200
Property taxes $2,700
Factory administration $16,810
Total $141,260
FLEXIBLE BUDGET PERFORMANCE REPORT
3
A B C D=B-C E=A-B
Panning Budget Flexible Budget Actual Cost Spending Variance Activity Variance
Direct labor $66,830 $63,570 $65,150 -$1,580 $3,260
Indirect labor $10,750 $10,450 $9,930 $520 $300
Utilities $8,780 $8,620 $9,170 -$550 $160
Supplies $1,710 $1,690 $1,920 -$230 $20
Equipment depreciation $29,780 $29,220 $29,220 $0 $560
Factory rent $8,200 $8,200 $8,600 -$400 $0
Property taxes $2,700 $2,700 $2,700 $0 $0
Factory administration $16,990 $16,810 $16,220 $590 $180
Total $145,740 $141,260 $142,910 -$1,650 $4,480